OpenStudy (anonymous):

A polygon has 170 diagonals. How many sides it will have ?

OpenStudy (nikvist):

\[d_n=170,\] \[\frac{n(n-3)}{2}=170,n(n-3)=340=20\cdot 17\] \[n=20\]